The equilibrium constant K, for the oxidation of sulfur dioxide into sulfur trioxide is 5.6x10¹ at 350°C. If the equilibrium partial pressures of SO, and SO, are 0.64 atm and 0.58 atm, respectively, what is the partial pressure of O, in the system?


2SO2(g) + O2(g) 250,(g)

1
Expert's answer
2022-06-03T02:43:03-0400

2SO2 + O2 = 2SO3

Kc = 5,6 * 10-1 = 0,56

P (SO2) = 0,64

P (SO3) = 0,58

P (O2) = ?


"K_c = \frac {[SO_3]²}{[SO_2]²[O_2]}"


"[O_2]= \cfrac {[SO_3]²}{\frac{[SO_2]²}{K_c}}"


"[O_2]= \cfrac {[0,58]²}{\frac{[0,64]²}{0,56}}= 0,46"


ANSWER: 0,46